OpenStudy (samsan9):

simplify the expression below write your answer only positive exponents. (a^2b^2c/a^5c^2)(a^3b/c^4

OpenStudy (samsan9):

\[(\frac{ a^2b^2c }{ a^5c^2 })\frac{a^3b }{ c^4 }\]

OpenStudy (anonymous):

\(\huge \frac{a^2b^2c}{a^5c^2}\times \frac{a^3b}{c^4}=\frac{a^5b^3c}{a^5c^6}\\\huge =\frac{(a^5c)b^3}{(a^5c)c^5}=\frac{b^3}{c^5}\)